I do not know how to write the equation to solve this problem.
A cough medicine contains 25% alcohol. How much liquid should a pharmacist add to 120 milliliter of the medicine so it contains only 20% alcohol?

Answer by stanbon(75887)

A cough medicine contains 25% alcohol. How much liquid should a pharmacist add to 120 milliliter of the medicine so it contains only 20% alcohol?
-----------------------------
Equation:
alcohol + alcohol = alcohol
---
I'm going to assume you are adding liquid that has on alcohol in it.
---
0.25*120 + 0*x = 0.20(x+120)
---
Multiply both sides by 100 to get rid of the decimals:
25*120 + 0 = 20(x+120)
25*120 = 20x + 20*120
20x = 5*120
x = 5*6
x = 30 milliliters (amt of liquid that should be added)
=============================================================
Cheers,
Stan H.